# CordViP: Correspondence-based Visuomotor Policy for Dexterous Manipulation in Real-World

### v1

@@ Abstract
Achieving human-level dexterity in robots is a key objective in the field of robotic manipulation. Recent advancements in 3D-based imitation learning have shown promising results, providing an effective pathway to achieve this goal. However, obtaining high-quality 3D representations presents two key problems: (1) the quality of point clouds captured by a single-view camera is significantly affected by factors such as camera resolution, positioning, and occlusions caused by the dexterous hand; (2) the global point clouds lack crucial contact information and spatial correspondences, which are necessary for fine-grained dexterous manipulation tasks. To eliminate these limitations, we propose CordViP, a novel framework that constructs and learns correspondences by leveraging the robust 6D pose estimation of objects and robot proprioception. Specifically, we first introduce the interaction-aware point clouds, which establish correspondences between the object and the hand. These point clouds are then used for our pre-training policy, where we also incorporate object-centric contact maps and hand-arm coordination information, effectively capturing both spatial and temporal dynamics. Our method demonstrates exceptional dexterous manipulation capabilities with an average success rate of 90% in four real-world tasks, surpassing other baselines by a large margin. Experimental results also highlight the superior generalization and robustness of CordViP to different objects, viewpoints, and scenarios. Code and videos are available on https://aureleopku.github.io/CordViP.
Figure 1: We propose CordViP, a correspondence-based visuomotor policy for dexterous manipulation in the real world. (a) Left: We present the interaction-aware point clouds, which demonstrate robustness to different viewpoints while establishing correspondences between the object and the hand. (b) Right: Our method achieves promising results across four real-world dexterous manipulation tasks, showcasing exceptional generalization capabilities.
@@ I Introduction
To comprehensively evaluate our proposed CordViP, we conduct extensive real-world experiments on four dexterous tasks: PickPlace, FlipCup, Assembly, and Artimanip. Comparative results demonstrate that our method not only exhibits superior effectiveness, but also achieves remarkable performance with a minor number of expert demonstrations, highlighting its capability to learn efficiently from limited data. Furthermore, we find that CordViP generalizes well to various environmental perturbations, including varying lighting conditions, unseen objects, diverse scenarios, and different camera viewpoints, outperforming other baselines by a large margin.
@@
    We develop a pipeline based on robust 6D pose estimation of objects and robot proprioception, which enables the real-time acquisition of complete 3D representations and semantic flow in real-world environments and can effectively address the challenges posed by occlusions during dexterous manipulation.
@@ II-B Imitation Learning
Imitation learning (IL) allows a robot to directly learn from experts. Behavioral Cloning (BC) is one of the simplest imitation learning algorithms, which treats the problem of learning behavior as a supervised learning task [37]. The modeling methods commonly used in traditional BC, such as MSE, discretization [29], and K-Means [16], have limitations when modeling complex action distributions. They fail to effectively capture the diversity and nuances of human behavior [35]. Over the past few years, diffusion models have emerged as a new modeling approach in BC, becoming powerful tools that enable robots to learn from demonstrations, handle uncertainty, and perform complex multi-step tasks with precision. From the early applications of DDPMs to the recent innovations in BESO [44], OCTO [49], and CrossFormers [10], these models have continually pushed the boundaries of what’s possible in robotic behavior generation. While traditional BC policies typically rely on 2D image-based representations [7, 67, 51, 60, 28], recent advancements have extended imitation learning to 3D visual representations [66, 4, 13, 52, 53, 30]. These 3D approaches provide a more comprehensive understanding of spatial relationships and 3D structures, further enhancing robotic behavior learning.
@@ III-B Interaction-aware Generation of 3D Point Clouds
Real-to-Sim for Digital Twin Generation. To achieve the goal of obtaining a complete and accurate static point cloud of the manipulated object, we aim to reconstruct the digital twin from a single-view image. Referring to the approach TripoSR [50], we implement a 3D generation technique, which utilizes its strong priors and broad understanding of visual concepts in the 3D world to generate 3D digital assets. To ensure the accuracy of point cloud flow tracking, we maintain consistency between the geometric and material properties of the reconstructed assets and their real-world counterparts. Subsequently, we uniformly sample points on the surface of the generated digital twin to obtain the initial 3D point clouds, providing a robust and accurate initial observation for both 3D spatial perception and pose tracking.
@@ IV-A Experiment Setup
Tasks. We evaluate our approach on four dexterous fine-grained manipulation tasks, as shown in Figure 5. The episode length of each task will be limited to a maximum of 380 steps and each task is evaluated with 20 trials. We now briefly describe our tasks:
@@
Figure 5: Visualization of four dexterous manipulation tasks, with the right side showing the end state.
@@
Baselines. We compare our method with six state-of-the-art imitation learning algorithms, i.e., three vision-based policies(BCRNN [31], ACT [67], DP [7]) and three 3D-based policies(BCRNN+3D, ACT+3D, DP3 [66]). BCRNN+3D and ACT+3D are variants of BCRNN and ACT, respectively, in which the image input is substituted with point clouds, encoded using PointNet. In fact, we also explore a more novel imitation learning framework, RISE [53]. However, we believe this method is not suitable for dexterous hand manipulation tasks, as it focuses solely on 3D observation and lacks crucial joint state information. Further details on the baseline choices and result analysis can be found in the Appendix -C. We emphasize that the resolution of both the image and depth for all 2D and 3D baseline methods is kept consistent, and the point clouds are synthesized from the RGBD camera.
@@ IV-B Effectiveness
The results of the effectiveness experiments are given in Table I. Our proposed CordViP maintains a completion rate of over 85% across four tasks, significantly outperforming the other baselines. BCRNN and its 3D variant perform poorly in all real-world tasks. While the dexterous hand can locate and reach the object, a substantial finger jitter is observed during the grasping and contact-rich manipulation phases. The image-based policies, ACT and DP, achieve good performance in the flipcup and pickplace tasks. However, they struggle with the assembly task. This is mainly due to the significant occlusions during manipulation, where these image-based policies fail to effectively leverage spatial information. Compared to ACT, ACT+3D shows superior results in all tasks, highlighting the crucial importance of geometric structure and spatial information in policy learning.
@@
Inference Speed. CordViP also attains an efficient inference speed, as it eliminates the need for the farthest point sampling of point clouds while utilizing compact 3D representations. We evaluate the inference speed of CordViP compared with DP3, CordViP reaches a maximum of 12.84 FPS, surpassing DP3’s 11.79 FPS. This highlights that our approach not only achieves enhanced performance but also maintains low computational overhead during inference.
TABLE I: Main results of four real-world tasks. Each experiment is evaluated with 20 trials.
@@ IV-C Efficiency
The number of expert demonstrations plays a crucial role in the performance of imitation learning. To access the learning efficiency, we train ACT, DP, DP3 and our proposed CordViP on four tasks using varying quantities of expert demonstrations. As illustrated in Figure 7, CordViP exhibits superior performance, achieving higher accuracy with fewer demonstrations. Remarkably, even with just 10 demonstrations, CordViP effectively establishes correspondences, extracts spatial and geometric features, and maintains a high success rate. Furthermore, when provided with a sufficient number of demonstrations, CordViP also demonstrates excellent manipulation capabilities.
@@ IV-D Generalization
Generalize to different lighting conditions. We set up three different lighting conditions. As shown in Table II, our proposed CordViP, based on 3D point clouds without color information, demonstrates a strong generalizability against various lighting conditions. In contrast, diffusion policy fails to complete tasks under most of these conditions. These image-based policies often rely on data augmentation to enhance generalization, which may lead to instability during training.
TABLE II: Generalization results on different lighting conditions. We evaluate the policy under three lighting scenarios: dim light(dim), white light(white) and colored lighting(colored).
@@
Generalize to different scenarios. CordViP relies on a robust 6D pose estimator to track objects and build interaction-aware 3D point clouds, which enables our method to generalize to different scenarios. The results presented in Table III show that DP is highly sensitive to visual variations, while ACT+3D performs well in scenarios with different visual appearances. However, 3D-based policy struggles in more challenging, cluttered scenes, as randomly placed objects can interfere with the synthesis of point clouds. In contrast, our method effectively focuses on the manipulation of subjects and objects, establishing spatial correspondences, and has demonstrated strong generalization to different scenarios.
TABLE III: Generalization results in diverse scenarios, including varying visual appearances and challenging cluttered environments.
@@
Generalize to unseen objects. To validate policies’ generalization to different objects, we test them on the PickPlace and FlipCup tasks using three unseen objects, each varying in color, shape, and dynamics. As shown in Table IV, DP shows poor generalization to unseen objects, while ACT+3D shows a certain level of generalization ability. However, ACT+3D struggles to adjust effectively and promptly when confronted with objects that exhibit significant differences. CordViP demonstrates strong generalization ability when handling unseen objects, making fine-grained adjustments by establishing spatial and temporal correspondences.
TABLE IV: Generalization results on unseen objects. For PickPlace and FlipCup tasks, we chose three previously unseen objects, varying in color, shape, and dynamics.
@@
Generalize to different viewpoints. Achieving generalization to different viewpoints presents a more significant challenge, as the expert demonstrations are all captured using a fixed camera. We evaluate DP, ACT+3D, and our method with three different camera viewpoints. For 3D-based methods, since the point clouds are represented in the world coordinate system, the camera is recalibrated for each viewpoint. As shown in Table V, the image-based diffusion policy is highly sensitive to camera viewpoints and completely fails across all three camera views. ACT+3D, which leverages 3D information, demonstrates a certain level of generalization to minor viewpoint changes. However, due to the limitations of the camera viewpoint, the synthesized point clouds can only capture partial spatial information, making it hard to handle significant changes in camera views. Our method, which generates interaction-aware point clouds, is capable of handling significant variations in camera perspectives.
TABLE V: Generalization results to different viewpoints.
@@
TABLE VI: Ablation experiments on the Effectiveness of different components. DP3 + Interaction-aware PC refers to using the Interaction-aware point clouds as the visual input for DP3. W/o. contact and coordination pretrain means that the encoder was not pre-trained with contact and coordination data.
@@ IV-E Ablations
Effectiveness of components. We conduct a series of ablation experiments on four real-world tasks to evaluate the effectiveness of different components in our method. As shown in Table VI, our results reveal the critical importance of contact and coordination information for learning dexterous manipulation policy. Specifically, for contact-rich tasks such as Flip-Cup, the absence of pre-training on contact information significantly reduces success rates. This highlights the essential role of contact in understanding physical interactions between the robotic hand and objects. On the other hand, for tasks that require continuous motion of both the arm and hand, such as Assembly and ArtiManip, incorporating coordination information notably enhances the policy’s ability to execute complex movement patterns. These findings emphasize the necessity of both contact and coordination information to achieve robust manipulation.
@@
Choice of point cloud encoders. We use PointNet [38], PointNet++ [39], and PointNeXt [41] as point cloud encoders. As shown in Table VII, PointNet outperforms the other encoders in success rates. We also freeze the point cloud encoder during the training phase of the correspondence-based diffusion policy. The results indicate that fine-tuning further enhances the learning of downstream tasks.
TABLE VII: Ablation experiments on point cloud encoders. * indicates that the encoder is frozen during the training phase of the correspondence-based diffusion policy.
@@ IV-F Transferability
Our proposed CordViP can be easily transferred to different backbones. Following Zhao et al. [67], we choose a Transformer-based CVAE as the decision backbone for the policy, which outputs a sequence of actions. We replace the 2D image input with our Interaction-aware point clouds and use the pre-trained encoder to establish correspondences between various components. Table VIII illustrates that CordViP achieves excellent performance on both diffusion-based and transformer-based backbones, showcasing strong transferability.
TABLE VIII: Transferability to different backbone.
@@ IV-G Failure Case Analysis
In this section, we analyze some failure cases of CordViP. We take the flipcup task as an example and visualize the results in Figure 8(a). We observe that while CordViP generally achieves good localization and grasping, the hand occasionally hovers in place after wrist rotation. One possible reason is that, during the collection of expert demonstrations, we deliberately reduced wrist speed, resulting in slight jitters that led the model to learn suboptimal solutions. A potential solution could be to expand the policy’s values of $horizon$ and $n\_action\_steps$.

### v2

@@ Abstract
Achieving human-level dexterity in robots is a key objective in the field of robotic manipulation. Recent advancements in 3D-based imitation learning have shown promising results, providing an effective pathway to achieve this goal. However, obtaining high-quality 3D representations presents two key problems: (1) the quality of point clouds captured by a single-view camera is significantly affected by factors such as camera resolution, positioning, and occlusions caused by the dexterous hand; (2) the global point clouds lack crucial contact information and spatial correspondences, which are necessary for fine-grained dexterous manipulation tasks. To eliminate these limitations, we propose CordViP, a novel framework that constructs and learns correspondences by leveraging the robust 6D pose estimation of objects and robot proprioception. Specifically, we first introduce the interaction-aware point clouds, which establish correspondences between the object and the hand. These point clouds are then used for our pre-training policy, where we also incorporate object-centric contact maps and hand-arm coordination information, effectively capturing both spatial and temporal dynamics. Our method demonstrates exceptional dexterous manipulation capabilities, achieving state-of-the-art performance in six real-world tasks, surpassing other baselines by a large margin. Experimental results also highlight the superior generalization and robustness of CordViP to different objects, viewpoints, and scenarios. Code and videos are available on https://aureleopku.github.io/CordViP.
Figure 1: We propose CordViP, a correspondence-based visuomotor policy for dexterous manipulation in the real world. (a) Left: We present the interaction-aware point clouds, which demonstrate robustness to different viewpoints while establishing correspondences between the object and the hand. (b) Right: Our method achieves promising results across multiple real-world dexterous manipulation tasks, showcasing exceptional generalization capabilities.
@@ I Introduction
To comprehensively evaluate our proposed CordViP, we conduct extensive real-world experiments on six dexterous tasks: PickPlace, FlipCup, Assembly, Artimanip, FlipCap and LongHoriManip. Comparative results demonstrate that our method not only exhibits superior effectiveness, but also achieves remarkable performance with a minor number of expert demonstrations, highlighting its capability to learn efficiently from limited data. Furthermore, we find that CordViP generalizes well to various environmental perturbations, including varying lighting conditions, unseen objects, diverse scenarios, and different camera viewpoints, outperforming other baselines by a large margin.
@@
    We develop a pipeline based on robust 6D pose estimation of objects and robot proprioception. This pipeline enables the real-time acquisition of complete 3D representations and semantic flow in real-world environments, and can effectively address the challenges posed by occlusions during dexterous manipulation.
@@ II-B Imitation Learning
Imitation learning (IL) allows a robot to directly learn from experts. Behavioral Cloning (BC) is one of the simplest imitation learning algorithms, which treats the problem of learning behavior as a supervised learning task [40]. The modeling methods commonly used in traditional BC, such as MSE, discretization [29], and K-Means [16], have limitations when modeling complex action distributions. They fail to effectively capture the diversity and nuances of human behavior [38]. Over the past few years, diffusion models have emerged as a new modeling approach in BC, becoming powerful tools that enable robots to learn from demonstrations, handle uncertainty, and perform complex multi-step tasks with precision. From the early applications of DDPMs to the recent innovations in BESO [47], OCTO [52], and CrossFormers [10], these models have continually pushed the boundaries of what’s possible in robotic behavior generation. While traditional BC policies typically rely on 2D image-based representations [7, 70, 54, 63, 28, 30], recent advancements have extended imitation learning to 3D visual representations [69, 4, 13, 55, 56, 31]. These 3D approaches provide a more comprehensive understanding of spatial relationships and 3D structures, further enhancing robotic behavior learning.
@@ III-B Interaction-aware Generation of 3D Point Clouds
Real-to-Sim for Digital Twin Generation. To achieve the goal of obtaining a complete and accurate static point cloud of the manipulated object, we aim to reconstruct the digital twin from a single-view image [34, 35]. Referring to the approach TripoSR [53], we implement a 3D generation technique, which utilizes its strong priors and broad understanding of visual concepts in the 3D world to generate 3D digital assets. To ensure the accuracy of point cloud flow tracking, we maintain consistency between the geometric and material properties of the reconstructed assets and their real-world counterparts. Subsequently, we uniformly sample points on the surface of the generated digital twin to obtain the initial 3D point clouds, providing a robust and accurate initial observation for both 3D spatial perception and pose tracking.
@@ IV-A Experiment Setup
Tasks. We evaluate our approach on four base dexterous manipulation tasks, along with two advanced contact-rich and fine-grained tasks, as shown in Figure 5. The episode length of each task will be limited to a maximum of 500 steps and each task is evaluated with 20 trials by default. We now briefly describe our tasks:
@@
  5. 5.
    FlipCap. This task requires four-finger coordination: the thumb and middle finger lift and rotate the cap slightly, while the index finger pushes it from the opposite side to complete the flip.
  6. 6.
    LongHoriManip. This task involves four sequential steps: pull, pick, place, and push, requiring precise and continuous control across a long horizon.
Figure 5: Visualization of six dexterous manipulation tasks, with the right side showing the end state.
@@
Baselines. We first compare our method with six state-of-the-art imitation learning algorithms, i.e., three vision-based policies(BCRNN [32], ACT [70], DP [7]) and three 3D-based policies(BCRNN+3D, ACT+3D, DP3 [69]). BCRNN+3D and ACT+3D are variants of BCRNN and ACT, respectively, in which the image input is substituted with point clouds, encoded using PointNet.We emphasize that the resolution of both the image and depth for all 2D and 3D baseline methods is kept consistent, and the point clouds are synthesized from the RGBD camera.
To further assess the performance under the same privileged setting as CordViP, we also introduce three additional baselines: (1) State-based MLP, a simple behavior cloning policy that takes the robot’s proprioception and the object pose as input; (2) State-based Diffusion Policy, which conditions a diffusion model on the concatenated proprioceptive and object pose features; and (3) G3Flow [4], a recent method that dynamically computes semantic flow based on pose information and integrates it with raw 3D point clouds in the Diffusion Policy framework.
@@ IV-B Effectiveness
The results of the effectiveness experiments are given in Table I and Table II. Our proposed CordViP maintains a completion rate of over 85% across four base tasks, significantly outperforming the other baselines, and also demonstrates superior performance on the two advanced tasks. BCRNN and its 3D variant perform poorly in all base tasks. While the dexterous hand can locate and reach the object, a substantial finger jitter is observed during the grasping and contact-rich manipulation phases. The image-based policies, ACT and DP, achieve good performance in the flipcup and pickplace tasks. However, they struggle with the assembly task. This is mainly due to the significant occlusions during manipulation, where these image-based policies fail to effectively leverage spatial information. state-based DP matches and even outperforms vision-based baselines on simpler tasks. However, due to the lack of geometric information, it struggles with more dexterous tasks. Compared to ACT, ACT+3D shows superior results in all tasks, highlighting the crucial importance of geometric structure and spatial information in policy learning. G3Flow, which benefits from both 3D visual and state information, is the strongest baseline overall. However, its performance on tasks such as FlipCap is limited due to the impact of 3D noise and the lack of correspondence modeling.
@@
Inference Efficiency. CordViP also attains an efficient inference speed, as it eliminates the need for the farthest point sampling of point clouds while utilizing compact 3D representations. We evaluate the inference speed of CordViP compared with DP3 on an Intel i7-14700KF CPU and RTX 4090D GPU, CordViP reaches a maximum of 12.84 FPS, surpassing DP3’s 11.79 FPS. This highlights that our approach not only achieves enhanced performance but also maintains low computational overhead during inference.
TABLE I: Main results of four base real-world tasks. Each experiment is evaluated with 20 trials.
@@
TABLE II: Main results of advanced tasks. Each experiment is evaluated with 10 trials.
Method FlipCap LongHoriManip
Pull Pick Place Push
ACT 30% 100% 40% 40% 30%
ACT+3D 40% 100% 40% 20% 20%
DP 30% 70% 30% 10% 10%
State-based DP 20% 80% 30% 10% 0%
G3Flow 60% 100% 80% 50% 30%
CordViP(Ours) 80% 100% 100% 70% 60%
@@ IV-C Efficiency
The number of expert demonstrations plays a crucial role in the performance of imitation learning. To access the learning efficiency, we train ACT, DP, DP3 and our proposed CordViP on dexterous tasks using varying quantities of expert demonstrations. As illustrated in Figure 7, CordViP exhibits superior performance, achieving higher accuracy with fewer demonstrations. Remarkably, even with just 10 demonstrations, CordViP effectively establishes correspondences, extracts spatial and geometric features, and maintains a high success rate. Furthermore, when provided with a sufficient number of demonstrations, CordViP also demonstrates excellent manipulation capabilities.
@@ IV-D Generalization
Generalize to different lighting conditions.
We established three distinct lighting conditions. As shown in Table III, the 3D representation of our proposed CordViP is dynamically maintained through FoundationPose tracking. FoundationPose, which integrates multimodal information (RGB images, depth data, and prior 3D model knowledge of the target object), demonstrated significant robustness to lighting variations in our empirical tests. In contrast, diffusion policies failed to complete tasks under most conditions. These image-based policies typically rely on data augmentation to improve generalization, which may introduce instability during training.
TABLE III: Generalization results on different lighting conditions. We evaluate the policy under three lighting scenarios: dim light(dim), white light(white) and colored lighting(colored).
@@
Generalize to different scenarios. CordViP relies on a robust 6D pose estimator to track objects and build interaction-aware 3D point clouds, which enables our method to generalize to different scenarios. The results presented in Table IV show that DP is highly sensitive to visual variations, while ACT+3D performs well in scenarios with different visual appearances due to its reliance on 3D geometric inputs. However, we observe 3D-based policy struggles in more challenging, cluttered scenes, as randomly placed objects can lead to noisy or incomplete point cloud representations. In contrast, our method effectively focuses on the manipulation of subjects and objects, establishing spatial correspondences, and has demonstrated strong generalization to different scenarios.
TABLE IV: Generalization results in diverse scenarios, including varying visual appearances and challenging cluttered environments.
@@
Generalize to unseen objects. To validate policies’ generalization to different objects, we test them on the PickPlace and FlipCup tasks using three unseen objects, each varying in color, shape, and dynamics. As shown in Table V, DP shows poor generalization to unseen objects, while ACT+3D shows a certain level of generalization ability. However, ACT+3D struggles to adjust effectively and promptly when confronted with objects that exhibit significant differences. CordViP demonstrates strong generalization ability when handling unseen objects, making fine-grained adjustments by establishing spatial and temporal correspondences.
TABLE V: Generalization results on unseen objects. For PickPlace and FlipCup tasks, we chose three previously unseen objects, varying in color, shape, and dynamics.
@@
Generalize to different viewpoints. Achieving generalization to different viewpoints presents a more significant challenge, as the expert demonstrations are all captured using a fixed camera. We evaluate DP, ACT+3D, and our method with three different camera viewpoints. For 3D-based methods, since the point clouds are represented in the world coordinate system, the camera is recalibrated for each viewpoint. As shown in Table VI, the image-based diffusion policy is highly sensitive to camera viewpoints and completely fails across all three camera views. ACT+3D, which leverages 3D information, demonstrates a certain level of generalization to minor viewpoint changes. However, due to the limitations of the camera viewpoint, the synthesized point clouds can only capture partial spatial information, making it hard to handle significant changes in camera views. By leveraging comprehensive 3D priors of manipulated objects, our method achieves view-agnostic generation of interaction-aware point clouds with full spatial coverage, demonstrating robust performance under significant variations in camera perspectives.
TABLE VI: Generalization results to different viewpoints.
@@
((a)) Case1: FlipCup.
((b)) Case2: Assembly.
TABLE VII: Ablation experiments on the Effectiveness of different components. DP3 + Interaction-aware PC refers to using the Interaction-aware point clouds as the visual input for DP3. W/o. contact and coordination pretrain means that the encoder was not pre-trained with contact and coordination data.
@@ IV-E Ablations
Effectiveness of components. We conduct a series of ablation experiments on four base real-world tasks to evaluate the effectiveness of different components in our method. As shown in Table VII, our results reveal the critical importance of contact and coordination information for learning dexterous manipulation policy. Specifically, for contact-rich tasks such as Flip-Cup, the absence of pre-training on contact information significantly reduces success rates. This highlights the essential role of contact in understanding physical interactions between the robotic hand and objects. On the other hand, for tasks that require continuous motion of both the arm and hand, such as Assembly and ArtiManip, incorporating coordination information notably enhances the policy’s ability to execute complex movement patterns. These findings emphasize the necessity of both contact and coordination information to achieve robust manipulation.
@@
Choice of point cloud encoders. We use PointNet [41], PointNet++ [42], and PointNeXt [44] as point cloud encoders. As shown in Table VIII, PointNet outperforms the other encoders in success rates. We also freeze the point cloud encoder during the training phase of the correspondence-based diffusion policy. The results indicate that fine-tuning further enhances the learning of downstream tasks.
TABLE VIII: Ablation experiments on point cloud encoders. * indicates that the encoder is frozen during the training phase of the correspondence-based diffusion policy.
@@ IV-F Transferability
Our proposed CordViP can be easily transferred to different backbones. Following the design of Zhao et al. [70], we choose a Transformer-based CVAE as one of the decision-making backbones for the policy, which outputs a sequence of actions. We replace the 2D image input with our Interaction-aware point clouds and use the pre-trained encoder to establish correspondences between various components. Table IX illustrates that CordViP achieves excellent performance on both diffusion-based and transformer-based backbones, showcasing strong transferability.
TABLE IX: Transferability to different backbone.
@@ IV-G Failure Case Analysis
In this section, we analyze some failure cases of CordViP. We take the flipcup task as an example and visualize the results in Figure 8(a). We observe that while CordViP generally achieves good localization and grasping, the hand occasionally hovers in place after wrist rotation. One possible reason is that, during the collection of expert demonstrations, we deliberately reduced wrist speed, resulting in slight jitters that led the model to learn suboptimal solutions. A potential solution could be expanding the policy’s planning horizon by increasing the values of $horizon$ and $n\_action\_steps$.